OpenStudy (anonymous):

A 25-mL solution of H2SO4 is completely neutralized by 18 mL of 1.0M NaOH. What is the concentration of the H2SO4? H2SO4(aq) +2NaOH(aq) → Na2SO4(aq) + 2H2O(l) A.0.72M B.0.009M C.0.50M D.0.36M

OpenStudy (anonymous):

We have M1V1=M2V2. Here M1 is to be found, V1=25ml=(25/1000)L & M2=1.0M, V2=18ml=(18/1000)L. Therefore M1=(M2V2)V1 \[M1=((18\div 1000)\times1.0)\div(25\div1000)=(18\div1000)\div(25\div1000)=(18\div1000)\times(1000\div25)=18\div25=0.72\] Hence, (A) 0.72M is the answer. :) Hope you get it.